# Jackknife Empirical Likelihood Method for U Statistics Based on Multivariate Samples and its Applications

### v1

@@ 6 Concluding remarks
To obtain expression for $EV_{l}$, we firstly simplify $V_{l}^{\prime}s$.
@@
$\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\hskip-7.22743pt=h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}-n_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}-n_{1}},\mathbb{Z}_{k_{1}-n_{1}-n_{2}}\ldots,\mathbb{Z}_{k_{m_{3}}-n_{1}-n_{2}}).$
@@
$=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg{[}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{i}\neq l,\;i=1,2,\dots,m_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}\hskip-14.45377ptI_{[1,n_{1}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
@@
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n-1\\ k_{k}\neq l,\;k=1,2,\dots,m_{3}\end{subarray}}I_{[n_{1}+n_{2}+1,n]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\Bigg{]},$ (19)
@@
Since $U_{n_{1},n_{2},n_{3}}=U_{n}$ and combining (2) and (5), (3) and (6), and (4) and (7), we have the following expressions,
@@
=($={n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{i,0,0}}{n_{1}-1}\right);$ (20)
@@
=($={n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,j,0}}{n_{2}-1}\right);$ (21)
@@
=($={n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,k}}{n_{3}-1}\right).$ (22)
@@
+($\qquad+{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,l-n_{1},0}}{n_{2}-1}\right)I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
+($\qquad+{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,l-n_{1}-n_{2}}}{n_{3}-1}\right)I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}$
@@
$\qquad+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}+\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}V_{l,0,0}I_{[1,n_{1}]}(l)$
$\qquad+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}V_{0,l-n_{1},0}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}V_{0,0,l-n_{1}-n_{2}}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}.$ (23)
Therefore the possible values of $EV_{l}$ are as follows
@@
$\qquad+\Bigg{[}\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}\theta I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}\theta I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}\theta I_{[n_{1}+n_{2}+1,n]}(l)\Bigg{]}.$
@@
To prove the Theorem 2.2, we first provide some lemmas, which play a vital role in the proof.
The following lemma ensures both the existence and uniqueness of the solution to equation (16). Its proof is analogous to the one given by [7] for the two-sample case, hence omitted.
@@
$S_{n}^{2}=nS_{n_{1},n_{2},n_{3}}^{2}+o_{p}(1)\;\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
@@
$\qquad\qquad\qquad\qquad+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg{]}\left(U_{n}-\theta\right)^{2}.$ (*)
@@
Moreover, the first term of equation (*) is (see Chapter 2 of [25])
@@
$\quad+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg{]}$
@@
The proof of the following Lemma is straight forward and on the same lines as of [1], and hence omitted.
@@
Noting that, (16) together with the fundamental inequality $\mid x\pm$ $y|\geq|x|-|y\mid$ leads to
00 0 $=\left|\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|=\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|$
@@
$|\lambda|=O_{p}\left(n^{-1/2}\right).$ (24)
@@
$\max_{1\leq l\leq n}\left|\gamma_{l}\right|$ $=|\lambda|\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=O_{p}\left(n^{-1/2}\right)o\left(n^{1/2}\right)=o_{p}(1).$ (25)
From (16), we have
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
@@
Also, using Lemma 6.3 and equations (24)-(25), last term in above expression is bounded by
@@
$-2\log R(\theta)=2\sum_{1=1}^{n}\gamma_{i}-\sum_{1=1}^{n}{\gamma_{i}}^{2}+\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}-\frac{1}{2}\sum_{1=1}^{n}{\gamma_{i}}^{4}+\ldots$ (26)
Let us first consider the third term of (26)
@@
Next, we consider the second term of (26) as
@@
Now, consider the first term of (26)
@@
$-2\log R(\theta)=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1)=\frac{(U_{n}-\theta)^{2}}{S_{n_{1},n_{2},n_{3}}^{2}}+o_{p}(1)\;\;\;\;\text{as }\;n_{1}\rightarrow\infty.$ (By using Lemma 6.2)
By using part (b) of Theorem 2.1 and the Slutsky’s theorem, $-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom. Hence the theorem proved.

### v2

@@ 6. Concluding remarks
We simplify the expression for $V_{l}^{\prime}s$ and then find $EV_{l}$.
@@
$\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\thinspace=h(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}-n_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}-n_{1}},\mathbb{Z}_{k_{1}-n_{1}-n_{2}}\ldots,\mathbb{Z}_{k_{m_{3}}-n_{1}-n_{2}}).$
@@
$=$ $nU_{n}-\frac{n(n-1)}{n-m}\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\Bigg[\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{i}\neq l,\;i=1,2,\dots,m_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\end{subarray}}\thinspace I_{[1,n_{1}]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}$
@@
$\qquad\qquad\qquad+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n-1\\ k_{k}\neq l,\;k=1,2,\dots,m_{3}\end{subarray}}I_{[n_{1}+n_{2}+1,n]}(l)\,\bar{h}_{i_{1},\dots,i_{m_{1}},j_{1},\dots,j_{m_{2}},k_{1},\dots,k_{m_{3}}}\Bigg],$ (6.1)
@@
Since $U_{n_{1},n_{2},n_{3}}=U_{n}$ and combining (2.2) and (2.5), (2.3) and (2.6), and (2.4) and (2.7), we have the following expressions,
@@
$={n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{1}U_{n}}{n_{1}-1}-\frac{V_{i,0,0}}{n_{1}-1}\right);$ (6.2)
@@
$={n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,j,0}}{n_{2}-1}\right);$ (6.3)
@@
$={n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,k}}{n_{3}-1}\right).$ (6.4)
@@
$\hskip 18.49988pt+{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}\left(\frac{n_{2}U_{n}}{n_{2}-1}-\frac{V_{0,l-n_{1},0}}{n_{2}-1}\right)I_{[n_{1}+1,n_{1}+n_{2}]}(l)$
$\hskip 18.49988pt+{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}\left(\frac{n_{3}U_{n}}{n_{3}-1}-\frac{V_{0,0,l-n_{1}-n_{2}}}{n_{3}-1}\right)I_{[n_{1}+n_{2}+1,n]}(l)\Bigg]$
@@
$\hskip 18.49988pt+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg]+\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}V_{l,0,0}I_{[1,n_{1}]}(l)$
$\hskip 18.49988pt+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}V_{0,l-n_{1},0}I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}V_{0,0,l-n_{1}-n_{2}}I_{[n_{1}+n_{2}+1,n]}(l)\Bigg].$ (6.5)
Therefore the values of $EV_{l}$ are as follows
@@
$\hskip 17.00024pt+\Bigg[\frac{n_{1}-m_{1}}{n_{1}(n_{1}-1)}\theta I_{[1,n_{1}]}(l)+\frac{n_{2}-m_{2}}{n_{2}(n_{2}-1)}\theta I_{[n_{1}+1,n_{1}+n_{2}]}(l)+\frac{n_{3}-m_{3}}{n_{3}(n_{3}-1)}\theta I_{[n_{1}+n_{2}+1,n]}(l)\Bigg].$
@@
We first provide some lemmas which play a vital role in the proof of Theorem 2.3.
The following lemma ensures both the existence and uniqueness of the solution to equation (2.16). Its proof is analogous to the one given by jing2009jackknife for the two-sample case, hence omitted.
@@
$S_{n}^{2}=n\sigma^{2}+o_{p}(1)\;\;\;\;\text{as }\;\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt+n_{3}\left(n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right)^{2}\bigg]\left(U_{n}-\theta\right)^{2}.$ (*)
@@
Moreover, the first term of equation (*) is (see Chapter 2 of wang2010empirical)
@@
$\qquad+\frac{(n_{2}-m_{2})^{2}}{n_{2}(n_{2}-1)}\frac{1}{n_{2}(n_{2}-1)}\sum_{l=n_{1}+1}^{n_{1}+n_{2}}\left(V_{0,l,0}-\bar{V}_{0,\cdot,0}\right)^{2}+\frac{(n_{3}-m_{3})^{2}}{n_{3}(n_{3}-1)}\frac{1}{n_{3}(n_{3}-1)}\left(V_{0,0,l}-\bar{V}_{0,0,\cdot}\right)^{2}\bigg]$
@@
The proof of the following Lemma is straight forward and on the same lines as of an2018jackknife, and hence omitted.
@@
Noting that, (2.16) together with the fundamental inequality $\mid x\pm$ $y|\geq|x|-|y\mid$ leads to
$0$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|=\frac{1}{n}\left|\sum_{i=1}^{n}\left(V_{i}-EV_{i}\right)-\lambda\sum_{i=1}^{n}\frac{\left(V_{i}-EV_{i}\right)^{2}}{1+\lambda\left(V_{i}-EV_{i}\right)}\right|$
@@
$|\lambda|=O_{p}\left(n^{-1/2}\right).$ (6.6)
@@
$\max_{1\leq l\leq n}\left|\gamma_{l}\right|$ $=|\lambda|\max_{1\leq l\leq n}\left|V_{l}-EV_{l}\right|=O_{p}\left(n^{-1/2}\right)o\left(n^{1/2}\right)=o_{p}(1).$ (6.7)
From (2.16), we have
$0$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}$
@@
Also, using Lemma 6.3 and equations (6.6)-(6.7), last term in the above expression is bounded by
@@
$2\log R(\theta)=2\sum_{1=1}^{n}\gamma_{i}-\sum_{1=1}^{n}{\gamma_{i}}^{2}+\frac{2}{3}\sum_{1=1}^{n}\gamma_{i}{}^{3}-\frac{1}{2}\sum_{1=1}^{n}{\gamma_{i}}^{4}+\ldots$ (6.8)
Let us first consider the third term of (6.8)
@@
Next, we consider the second term of (6.8) as
@@
Now, consider the first term of (6.8)
@@
$-2\log R(\theta)=\frac{n(U_{n}-\theta)^{2}}{S_{n}^{2}}+o_{p}(1)=\frac{(U_{n}-\theta)^{2}}{\sigma^{2}}+o_{p}(1)\;\;\;\;\text{as }\;n_{1}\rightarrow\infty.$ (By using Lemma 6.2)
By using part (b) of Theorem 2.1 and the Slutsky’s theorem, $2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom. Hence the theorem is proved.